# Corresponding Author

### v1

@@ Abstract
Multimodal Large Language Models (MLLMs) have emerged as a promising way to automate Radiology Report Generation (RRG). In this work, we systematically investigate the design space of 3D MLLMs, including visual input representation, projectors, Large Language Models (LLMs), and fine-tuning techniques for 3D CT report generation. We also introduce two knowledge-based report augmentation methods that improve performance on the GREEN score by up to 10%, achieving the 2nd place on the MICCAI 2024 AMOS-MM challenge. Our results on the 1,687 cases from the AMOS-MM dataset show that RRG is largely independent of the size of LLM under the same training protocol. We also show that larger volume size does not always improve performance if the original ViT was pre-trained on a smaller volume size. Lastly, we show that using a segmentation mask along with the CT volume improves performance. The code is publicly available at https://github.com/bowang-lab/AMOS-MM-Solution.
@@ 1 Introduction
Computed Tomography (CT) is a cornerstone of modern diagnostic imaging, offering detailed insights into internal anatomical structures and playing a critical role in diagnosing a wide range of diseases [7]. However, the rapid increase in the need of CT examinations presents a significant challenge for radiologists, who must interpret complex 3D volumetric data and generate comprehensive reports under tight time constraints [4]. This growing demand places immense pressure on healthcare systems, often leading to delays in diagnosis and treatment, which can adversely affect patient outcomes [24].
@@ 2.4 Knowledge-base Report Augmentation
We further explore ways to augment the generated reports to ensure completeness, using two introduced methods: Binary-based Questioning (BQ) and naive normality (NN) augmentation. Fig 2 shows a graphical representations of the two methods.
Figure 2: Knowledge-base report augmentation. After the MLLM generates a report, we apply two additional methods to augment the report, which are Binary-based Questioning (BQ) and naive normality. For BQ, we train a triplet model to answer questions about common conditions and use its answers to append additional findings to the generated report. To train the triplet model, we first prepare a set of question-answer pairs by prompting GPT o3-mini to generate triplets of the entity, location, and existence (in the format {entity, location, exsist}) using findings from the reference reports. These triplets are then reformatted to questions, which are used to prompt the triplet model to generate binary True/False answers based on the CT volume. The value of "exist" is used as ground truth. At inference time, we prompt our triplet model with common triplets, which correspond to common findings, and based on its binary answer we append a positive or negative finding to the report. For naive normality, we append normal findings for organs that are not mentioned in the generated report after the BQ step. These two methods go hand-in-hand to ensure that our reports are complete and do not miss any common findings.
@@ 3.3 Results
We report medical report generation results for different LLMs, projectors, and fine-tuning methods. We also explore different settings such as volume size, predicting impressions along with findings, and adding segmentation masks. We use the GREEN score [21] as our base evaluation metric, but we also report RaTEScore [30] and text similarity metrics such as BLEU [22], ROGUE [16] and METEOR [12].
@@
We further experiment with several different training scenarios, such as predicting impressions along with findings and feeding a segmentation mask generated by TotalSegmentor [25], along with the CT volume, where we embed the volume and organ segmentation mask separately and concatenate their embeddings before being passed to the LLM. The results are shown in Table 3. Predicting impressions slightly degrades performance. However, using the segmentation mask along with the volume slightly improves performance on both GREEN and RaTEScore, which could be because it’s easier to identify abnormalities and organs with the help of the masks.
@@
Figure 3: Qualitative examples of augmented reports. We show 3 examples of ground truth, generated, and the augmented reports after applying Binary-based Questioning and Naive Normality. Findings highlighted in red represents findings that are originally missed in the generated report, but that are then captured in the augmented one.Bolded findings in the augmented report represents the original generated report. All three examples show that our report augmentation methods are able to capture missed positive findings, while also ensuring completeness by mentioning normal organs explicitly.

### v2

@@ Abstract
Multimodal Large Language Models (MLLMs) have emerged as a promising way to automate Radiology Report Generation (RRG). In this work, we systematically investigate the design space of 3D MLLMs, including visual input representation, projectors, Large Language Models (LLMs), and fine-tuning techniques for 3D CT report generation. We also introduce two knowledge-based report augmentation methods that improve performance on the GREEN score by up to 10%, achieving the 2nd place on the MICCAI 2024 AMOS-MM challenge. Our results on the 1,687 cases from the AMOS-MM dataset show that RRG is largely independent of the size of the LLM under the same training protocol. We also show that larger volume size does not always improve performance if the original ViT was pre-trained on a smaller volume size. Lastly, we show that using a segmentation mask along with the CT volume improves performance. The code is publicly available at https://github.com/bowang-lab/AMOS-MM-Solution.
@@ 1 Introduction
Computed Tomography (CT) is a cornerstone of modern diagnostic imaging, offering detailed insights into internal anatomical structures and playing a critical role in diagnosing a wide range of diseases [7]. However, the rapid increase in the need for CT examinations presents a significant challenge for radiologists, who must interpret complex 3D volumetric data and generate comprehensive reports under tight time constraints [4]. This growing demand places immense pressure on healthcare systems, often leading to delays in diagnosis and treatment, which can adversely affect patient outcomes [24].
@@ 2.4 Knowledge-base Report Augmentation
We further explore ways to augment the generated reports to ensure completeness, using two introduced methods: Binary-based Questioning (BQ) and naive normality (NN) augmentation. Fig 2 shows a graphical representation of the two methods.
Figure 2: Knowledge-base report augmentation. After the MLLM generates a report, we apply two additional methods to augment the report, which are Binary-based Questioning (BQ) and naive normality. For BQ, we train a triplet model to answer questions about common conditions and use its answers to append additional findings to the generated report. To train the triplet model, we first prepare a set of question-answer pairs by prompting GPT o3-mini to generate triplets of the entity, location, and existence (in the format {entity, location, exist}) using findings from the reference reports. These triplets are then reformatted to questions, which are used to prompt the triplet model to generate binary True/False answers based on the CT volume. The value of "exist" is used as ground truth. At inference time, we prompt our triplet model with common triplets, which correspond to common findings, and based on its binary answer we append a positive or negative finding to the report. For naive normality, we append normal findings for organs that are not mentioned in the generated report after the BQ step. These two methods go hand-in-hand to ensure that our reports are complete and do not miss any common findings.
@@ 3.3 Results
We report medical report generation results for different LLMs, projectors, and fine-tuning methods. We also explore different settings such as volume size, predicting impressions along with findings, and adding segmentation masks. We use the GREEN score [21] as our base evaluation metric, but we also report RaTEScore [30] and text similarity metrics such as BLEU [22], ROUGE [16] and METEOR [12].
@@
We further experiment with several different training scenarios, such as predicting impressions along with findings and feeding a segmentation mask generated by TotalSegmentor [25], along with the CT volume, where we embed the volume and organ segmentation mask separately and concatenate their embeddings before being passed to the LLM. The results are shown in Table 3. Predicting impressions slightly degrades performance. However, using the segmentation mask along with the volume slightly improves performance on both GREEN and RaTEScore, which could be because it is easier to identify abnormalities and organs with the help of the masks.
@@
Figure 3: Qualitative examples of augmented reports. We show 3 examples of ground truth, generated, and the augmented reports after applying Binary-based Questioning and Naive Normality. Findings highlighted in red represents findings that were originally missed in the generated report, but that are then captured in the augmented one.Bolded findings in the augmented report represents the original generated report. All three examples show that our report augmentation methods are able to capture missed positive findings, while also ensuring completeness by mentioning normal organs explicitly.